OpenStudy (anonymous):

Word Problems Help! 12 men take 6 hours to finish a work. After the 12 men have worked for 1 hour, the contractor decides to call in 8 more men so that the work can be completed earlier. How many hrs would 20 men take to complete the remaining work?

hero (hero):

Just a minute please,

hero (hero):

@JustinSpeedster you still here?

OpenStudy (anonymous):

work = men/hour \[\frac{ 12 }{ 6 }\]\[\frac{ 12 }{ 6 }\ - \frac{ 1 }{ 5 }\] is that right?

OpenStudy (anonymous):

@Hero sorry for the afk...can you teach me how to solve it? I'm stuck.

hero (hero):

The key to solving this Is Keeping track of the total man hours

hero (hero):

It takes 12 men 6 hours to complete the Work

hero (hero):

multiplying 12 by 6 gets you 72 total man hours necessary to complete the job.

hero (hero):

They tell us that the 12 men have already worked 1 hour or 12 man ho urs so We Subtract 12 from 72 to get 60 Man hours remaining

hero (hero):

They add eight more men to this project so at that point , we have 20 men and 60 Man hours left to complete the job. How many hours will it take?

OpenStudy (anonymous):

I don't get it T.T

OpenStudy (anonymous):

5 hours?

hero (hero):

At this point you divide 60 by 20 to get the number of hours it will take the 20 men to complete the job.

OpenStudy (anonymous):

3 hours

hero (hero):

yes three hours, now try to understand the reasoning behind this approach.

OpenStudy (anonymous):

why did you multiply 12 by 6?

hero (hero):

To Calculate the total man hours necessary to complete the work.

OpenStudy (anonymous):

OH I see..I guess it was easy after analyzing it. I feel like a stupid now. The conflict that I encountered was the forming of equation. I failed to calculate the man hours....I was focused on solving the rate on how 1 worker/hour to solve everything. Oh well. Thanks for the help, HERO. Another lesson well taught. :D

hero (hero):

I don't like solving it using fractions. That approach is slightly unintuitive.

hero (hero):

my first train of thought is to calculate total man hours Instead of rate